find the equation of the line passing through the given point, write the equation in standard form AX+By=C (-7,-4) and (-1,-9)
Answer by mananth(16949)

x1 y1 x2 y2
-7 -4 -1 -9

slope m = (y2-y1)/(x2-x1)
( -9 - -4 )/( -1 - -7 )
( -5 / 6 )
m= - 5/ 6

Plug value of the slope and point (-7 ,-4) in
Y = m x + b
-4.00 = 35/6 + b
b= -4 - 35/6
b= -59/ 6
So the equation will be
Y = - 5/6 x -59/6
m.ananth@hotmail.ca